### v1

# Discrete homotopy and homology theories for finite posets

Abstract.
This paper presents a discrete homotopy theory and a discrete homology theory for finite posets. In particular, the discrete and classical homotopy groups of finite posets are always isomorphic. Moreover, this discrete homology theory is related to the discrete homotopy theory through a discrete analogue of the Hurewicz map.
Key words and phrases:
finite poset, discrete homotopy theory, discrete cubical homology theory, Hurewicz map
2020 Mathematics Subject Classification:
55N35, 55Q05, 06A07
1 Both authors are corresponding authors
1. Introduction
The classical homotopy and homology theories for finite posets are based on a well-known correspondence given by McCord, which assigns to each finite poset $M$ the order complex $\mathcal{K}(M)$ together with a weak homotopy homotopy equivalence $|\mathcal{K}(M)|\rightarrow M$ [1, Theorem 1.4.6]. This correspondence permits the study of homotopy and homology groups of a finite poset by means of its associated order complex.
The goal of this paper is to develop a discrete homotopy theory and a discrete homology theory, which are discrete analogues of homotopy and homology theories, associating a sequence of groups to a finite poset, capturing its combinatorial structure, rather than its topological structure. Furthermore, discrete homology theory is related to discrete homotopy theory in the same way that classical homology is related to classical homotopy theory. Since every finite digraph can be viewed as the Hasse diagram of a finite poset, our theories provide much algebraic insight into problems regarding the combinatorics of finite digraphs.
We prove three main results. The first is motivated by a discrete homotopy theory for graphs introduced in [2, 5, 8]. In similar fashion we define a discrete homotopy theory for finite posets. Our first result asserts that, for all finite posets, the discrete homotopy theory and the classical one lead to isomorphic homotopy groups in all dimensions (Theorem 3.3). This allows us to study the classical homotopy groups of compact polyhedra from a new viewpoint, the weak homotopy types of compact polyhedra being in one to one correspondence with the weak homotopy types of finite posets [1, Theorem 1.4.6, Theorem 1.4.12].
In general, the classical homotopy groups of topological spaces are surprisingly difficult to compute. For example, covering spaces are used to calculate the fundamental group of a circle, one of the most foundational calculations in algebraic topology. To illustrate how the methods of calculating discrete homotopy groups of some topological spaces can be simpler and more tractable than the ones used to get the classical homotopy groups, we give a new method for calculating the fundamental group of a circle by means of this new homotopy theory (Example 3.4).
The motivation for the second result comes from a discrete cubical homology theory for graphs introduced in [3, 4]. We define a similar discrete cubical homology theory for finite posets. Our second result is that we exhibit for each $n\geq 1$ a natural homomorphism $\psi:H_{n}^{\rm Cube}(M)\rightarrow H_{n}^{\rm Simpl}(\mathcal{K}(M))$, where $H_{n}^{\rm Cube}(M)$ is the $n$th discrete cubical homology group, and a class of finite posets $M$ for which $\psi$ is surjective (Theorem 4.7). We provide an example for which the homology groups of the two theories are not isomorphic at least in dimension $1$ (Example 4.9).
Theorem 4.7 has an interesting application as follows. Let $M=|K|$ be a triangulated $m$-manifold with $K$ simplicially collapsible. Corollary 4.17 tells when the manifold $M$ with two specific balls removed is not contractible.
The third result makes progress towards the discrete analogue of the Hurewicz map relating discrete homology theory to discrete homotopy theory, motivated by the one defined in [8]. Similarly, we give a notion of discrete Hurewicz map. There is a discrete Hurewicz theorem in dimension 1 (Theorem 5.4). Continuing our analogy with classical homotopy and homology theories, our third result asserts that this discrete Hurewicz map coincides with the classical one (Theorem 5.5).
The organization of the paper is as follows. In Section 2, we give preliminary definitions and results concerning finite posets and simplicial complexes. In Section 3, we define a discrete homotopy theory for finite posets and show that the discrete and classical homotopy groups of finite posets agree in all dimensions. In Section 4, we define a discrete cubical homology theory for finite posets, and explore both the similarities and differences between the discrete cubical homology theory and the classical one. In Section 5, we define a discrete Hurewicz map which is the same as the classical one.
In this paper, all maps between posets that we deal with are assumed to be order-preserving.
2. Preliminaries
This section is devoted to introducing some definitions and properties of finite posets and simplicial complexes that are needed in the sequel. We refer the reader to [1, 7, 11, 12] for further details or proofs.
2.1. Finite posets
A topological space is said to be Alexandroff if every intersection of open sets is still an open set. If $X$ is an Alexandroff space, for each $x\in X,$ we denote by $U_{x}$ the intersection of all the open sets of $X$ containing $x$. This allows us to define a reflexive and transitive relation on $X$ by declaring $x\leq y$ if $U_{x}\subseteq U_{y}.$ The relation is antisymmetric if and only if $X$ is $T_{0}$. On the other hand, if $X$ is a set endowed with a reflexive, transitive relation, then the sets $U_{x}=\{y\in X|y\leq x\}$, as a basis, determines an Alexandroff topology. The Alenxandroff topology is $T_{0}$ if and only if the relation is antisymmetric, that is, a partial order. A set with a partial order is called a poset. It has been shown that Alexandroff $T_{0}$ spaces and posets are in bijective correspondence. A finite $T_{0}$ space is clearly an Alexandroff $T_{0}$ space, thus a finite $T_{0}$ space is equivalent to a finite poset.
A useful way to represent posets is with their Hasse diagrams. The Hasse diagram of a poset $X$ is a digraph whose vertices are the points of $X$ and whose edges are the ordered pairs $(x,y)$ such that $x<y$ and there exists no $z\in X$ such that $x<z<y$.
Given a poset $X$ and an element $x\in X$, we denote $U_{x}=\{y\in X~{}|~{}y\leq x\},~{}\widehat{U}_{x}=\{y\in X~{}|~{}y<x\},$ $F_{x}=\{y\in X~{}|~{}x\leq y\},$ and $\widehat{F}_{x}=\{y\in X~{}|~{}x<y\}.$ Let $C_{x}=U_{x}\cup F_{x}$ and $\widehat{C}_{x}=\widehat{U}_{x}\cup\widehat{F}_{x}$. A totally ordered subset $C$ of $X$ is called a chain. An $n$-chain of $X$ is a chain of cardinality of $n+1$.
Definition 2.1.
A finite poset $X$ is homogeneous of dimension $n$ if all maximal chains in $X$ have cardinality of $n+1$. In this case, the degree of $x$, denoted ${\rm deg}(x)$, is the dimension of $U_{x}$.
If $X$ is a finite poset, we define its order complex $\mathcal{K}(X)$ as the ordered simplicial complex whose simplices are the nonempty chains of $X$. Conversely, given a finite simplicial complex $K$, we define the face poset $\chi(K)$ of $K$ as the poset of simplices of $K$ ordered by inclusion.
Definition 2.2.
Let $X$ be a finite poset. Define the K-McCord map $\mu_{X}:|\mathcal{K}(X)|\rightarrow X$ by $\mu_{X}(u)=\text{min}(\text{support}(u))$.
A result of McCord [1, Theorem 1.4.6] asserts that $\mu_{X}$ is a weak homotopy equivalence. In fact, the same argument in [1, Theorem 1.4.6] shows $\widehat{\mu_{X}}:|\mathcal{K}(X)|\rightarrow X$ defined by
(2.1) $\widehat{\mu_{X}}(u)=\text{max}(\text{support}(u))$
is also a weak homotopy equivalence.
Let $f,g:X\rightarrow Y$ be two maps between finite posets. The following statements hold.
Proposition 2.3.
  • (1)
    $f$ is continuous if and only if it is order-preserving.
  • (2)
    $f$ is homotopic to $g$ if and only if there is a sequence $f_{0},f_{1},\cdots,f_{n}$ of order-preserving maps such that $f=f_{0}\leq f_{1}\geq f_{2}\leq\cdots f_{n}=g$.
2.2. Simplicial complexes
Definition 2.4.
Let $K$ be a simplicial complex. A simplex $A\subseteq K$ is called a principle simplex if $A$ is not a proper face of any simplex of $K$. The face $B$ of $A$ is called a free face of $A$ in $K$ if $B$ is the proper face of no other simplex of $K$.
$K$ is a principle $k$-simplicial complex if every principle simplex of $K$ is $k$-dimensional.
Definition 2.5.
[…]
Definition 2.6.
Suppose $X\subseteq Y$ are polyhedra and that $Y=X\cup B^{n}$ and $X\cap B^{n}=$ a face $B^{n-1}$, where $B^{i}$ denotes a piecewise-linear (PL) $i$-ball. Then we say that there is an elementary collapse of $Y$ onto $X$, and write $Y\searrow X$. The collapse is across $B^{n}$ onto $B^{n-1}$ from the complementary face ${\rm cl}(\partial B^{n}-B^{n-1})$.
We say $Y$ collapses to $X$, writen $Y\searrow X$, if there is a finite sequence of elementary collapses $Y=Y_{0}\searrow Y_{1}\searrow\cdots\searrow Y_{n}=X$. If $X$ is a point, we call $Y$ collapsible and write $Y\searrow 0$.
[…]
3. Discrete homotopy theory for fintie posets
In this section we define a discrete homotopy theory for finite posets, motivated by the notion of a discrete homotopy theory for graphs introduced in [5, 8]. As we shall see in Theorem 3.3, the discrete and classical homotopy groups of finite posets agree in all dimensions.
Let $\mathbb{Z}$ denote the digraph whose vertices are integers and whose edges are the ordered pairs $(2i,2i+1)$ and $(2j,2j-1)$. Given an integer $p\geq 0$, write $I_{p}$ for the subgraph of $\mathbb{Z}$ induced by $\{0,\cdots,p\}$.
Two maps $f,g:X\rightarrow Y$ between finite posets are homotopic if for some $p$, there is a map $H:X\times I_{p}\rightarrow Y$ such that $H(x,0)=f(x)$ and $H(x,p)=g(x)$. The map $H$ is a homotopy from $f$ to $g$. Write $H_{i}(x)=H(x,i)$. This definition in fact coincides with Proposition 2.3.
Let $\mathbb{Z}^{n}$ be the $n$-fold Cartesian product of $\mathbb{Z}$ whose vertices are $n$-tuples of integers. For any $r\geq 0$, $\mathbb{Z}^{n}_{\geq r}$ denotes the subgraph consisting of all vertices $x=(x_{1},\cdots,x_{n})\in\mathbb{Z}^{n}$ with $|x_{i}|=r$ for some $i$, and $\mathbb{Z}^{n}_{\leq r}$ denotes the subgraph consisting of all vertices $x=(x_{1},\cdots,x_{n})\in\mathbb{Z}^{n}$ with $|x_{i}|\leq r$ for every $i$. Denote by $\partial\mathbb{Z}^{n}_{\leq r}$ the subgraph of $\mathbb{Z}^{n}_{\leq r}$ consisting of all vertices $x=(x_{1},\cdots,x_{n})\in\mathbb{Z}^{n}_{\leq r}$ with $|x_{i}|=r$ for some $i$. Similarly, Denote by $\partial I^{n}_{\leq r}$ the subgraph of $I^{n}_{\leq r}$ consisting of all vertices $x=(x_{1},\cdots,x_{n})\in I^{n}_{\leq r}$ with $x_{i}=r$ or 00 for some $i$.
By abuse of notation, we will write
$f:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\rightarrow(X,x_{0})$
if there exists an $r\geq 0$ such that $f$ is a map $f:(\mathbb{Z}^{n},\mathbb{Z}_{\geq r}^{n})\rightarrow(X,x_{0})$. The minimum of such $r$ is called the radius of $f$
Figure 1. The Hasse diagram of $\mathbb{Z}^{2}$.
For a finite poset $X$ with basepoint $x_{0}\in X$, two maps $f,g:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\rightarrow(X,x_{0})$ are homotopic if there is a basepoint-preserving homotopy $H:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\times I_{p}\rightarrow X$ from $f$ to $g$ such that $H_{i}$ is a map $(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\rightarrow(X,x_{0})$ for all $i$. The basepoint-preserving homotopy class of $f$ is denoted $[f]$.
Definition 3.1.
For a finite poset $X$ with basepoint $x_{0}\in X$, define $\pi_{n}^{D}(X,x_{0})$ to be the set of basepoint-preserving-homotopy classes of maps $f:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\rightarrow(X,x_{0})$. We call $\pi_{n}^{D}(X,x_{0})$ the $n$th discrete homotopy group of $X$.
We can endow $\pi^{D}_{n}(X,x_{0})$ with a group structure as follows. If $f$ and $g$ are maps $(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\rightarrow(X,x_{0})$ of radii $r_{f}$ and $r_{g}$ respectively, then the product $f\cdot g$ is defined by
$(f\cdot g)(x_{1},\cdots,x_{n})=\begin{cases}f(x_{1},\cdots,x_{n})~{}&\text{if}~{}x_{1}\leq r_{f},\\ g(x_{1}-(r_{f}+r_{g}),x_{2},\cdots,x_{n})~{}&\text{if}~{}x_{1}>r_{f}.\end{cases}$
The identity element is the constant map sending $\mathbb{Z}^{n}$ to $x_{0}$.
Similar arguments as for discrete homotopy groups of graphs in [5] show that the definition of $\pi^{D}_{n}(X,x_{0})$ is independent of the basepoint $x_{0}$ when $X$ is connected, and that $\pi^{D}_{n}(X,x_{0})$ is a group under the product defined above. Furthermore, the idea of the proof of [5, Proposition 3.6] can be used to show that $\pi^{D}_{n}(X,x_{0})$ is abelian for $n\geq 2$.
Before we construct a homomorphism from the discrete $n$-th homotopy group $\pi^{D}_{n}(X,x_{0})$ to the classical $n$-th homotopy group $\pi_{n}(|\mathcal{K}(X)|,x_{0})$, we make a few preliminary observations.
Let $L^{n}$ be the $n$-dimensional cube, the product of $n$ copies of the interval $[-1,1]$. Given a partition
(3.1) $-1=t_{-\widehat{m_{i}}}<t_{-\widehat{m_{i}}+1}<\cdots<t_{-1}<0=t_{0}<t_{1}\cdots<t_{m_{i}-1}<t_{m_{i}}=1,$
of the $i$-th copy of $L^{n}$ for each $i$, $\widehat{m_{i}}$ being a positive integer, $L^{n}$ can be identified with the geometric realization of a subgraph $N$ of $\mathbb{Z}^{n}$,
$N=\prod_{i}^{n}\mathbb{Z}_{[-\widehat{m_{i}},m_{i}]}.$
Write $P$ for this partition. The subgraph $N$ is denoted $\gamma_{P}$. For each vertex $(x_{1},x_{2},\cdots,x_{n})$ of $\gamma_{P}$, taking it to the point $(t_{x_{1}},t_{x_{2}},\cdots,t_{x_{n}})$ of $L^{n}$ induces a canonical homeomorphism
$\lambda_{P}:|\mathcal{K}(\gamma_{P})|$ $\longrightarrow L^{n}.$
When the partition $P$ takes the form
$-1=\frac{-r}{r}<\frac{-r+1}{r}<\cdots<\frac{-1}{r}<0=t_{0}<\frac{1}{r}\cdots<\frac{r-1}{r}<\frac{r}{r}=1$
for each $i$, we will write $P$, $\gamma_{P}$ and $\lambda_{P}$ as $P_{r}$, $\gamma_{r}$ and $\lambda_{r}$, respectively.
Let $I$ be the unit interval $[0,1]$. The geometric realization of the subgraph, $I_{q}$, of $\mathbb{Z}$ can be viewed as $I$ by the canonical homeomorphism
$\beta_{q}:|\mathcal{K}(I_{q})|\rightarrow I$
with $j$ mapping to $\frac{j}{q}$ for each vertex $j$.
Definition 3.2.
Let $X$ be a connected finite poset. Define $\theta:\pi_{n}^{D}(X,x_{0})\rightarrow\pi_{n}(|\mathcal{K}(X)|,x_{0})$ by the equation
$\theta([f])=[|\mathcal{K}(f)|\circ\lambda_{r_{f}}^{-1}],$
where $r_{f}$ is the radius of the map $f$.
It is straightforward to check that $\theta$ is well-defined and is a homomorphism.
Theorem 3.3.
The homomorphism $\theta$ is an isomorphism.
Proof.
Since $X$ is connected, we may assume that $x_{0}$ is a minimum vertex of $X$. Suppose $f:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\rightarrow(X,x_{0})$ is a map of radius $r$ such that $\theta([f])=0$. Let $G:(L^{n},\partial L^{n})\times I\rightarrow(|\mathcal{K}(X)|,x_{0})$ be a homotopy from $|\mathcal{K}(f)|\circ\lambda_{r}^{-1}$ to the constant map. Note that $G$ can be written as
$G:\left(\lambda_{r}(|\mathcal{K}(\gamma_{r})|),\partial\lambda_{r}(|\mathcal{K}(\gamma_{r})|)\right)\times|\mathcal{K}(I_{q})|\rightarrow(|\mathcal{K}(X)|,x_{0}).$
Cover $\lambda_{r}(|\mathcal{K}(\gamma_{r})|)\times|\mathcal{K}(I_{q})|$ by the open sets
$(\widehat{\mu_{X}}\circ G)^{-1}(U^{X}_{u}),$
as $u$ ranges over all vertices of $X$. Denote this open covers by $\mathcal{A}$.
Let $\delta_{\mathcal{A}}$ be the Lebesgue number for the open cover $\mathcal{A}$. We take $r_{1}$ to be a multiple of $r$ and $q_{1}$ to be a even integer larger than $q$ so that each subpolyhedron of $\lambda_{r_{1}}(|\mathcal{K}(\gamma_{r_{1}})|)$, which is the image of a simplex, has diameter less than $\frac{\delta}{2}$, where $\delta<\delta_{\mathcal{A}}$. Let us construct a homotopy
$H:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\times I_{q_{1}}\rightarrow(X,x_{0})$
as follows. If $(w,t)\in\mathbb{Z}_{\leq r_{1}}^{n}\times I_{q_{1}}-\mathbb{Z}_{\leq r_{1}}^{n}\times 0-\mathbb{Z}_{\leq r_{1}}^{n}\times q_{1}$, then we have
(3.2) $\lambda_{r_{1}}(|\mathcal{K}(U^{\gamma_{r_{1}}}_{w})|)\times\beta_{q_{1}}(|\mathcal{K}(U^{I_{q_{1}}}_{t})|)\subseteq(\widehat{\mu_{X}}\circ G)^{-1}(U^{X}_{u}),$
for some vertex $u$ of $X$, $\lambda_{r_{1}}(|\mathcal{K}(U^{\gamma_{r_{1}}}_{w})|)\times|\mathcal{K}(U^{I_{q_{1}}}_{t})|$ being a subset of the closed star of $(\lambda_{r_{1}}(w),t)$. In this case, set
$H(w,t)=u,$
[…]
(3.3) $H(w,t)=\begin{cases}x_{0}~{}&\text{if}~{}w\in\partial\mathbb{Z}_{\leq r_{1}}^{n}~{}\text{or}~{}t=q_{1},\\ \widehat{\mu_{X}}\circ(|\mathcal{K}(f)|\circ\lambda_{r}^{-1})\circ\lambda_{r_{1}}(w)~{}&\text{if}~{}(w,t)\in\mathbb{Z}_{\leq r_{1}}^{n}\times 0.\end{cases}$
Let us check that $H$ is an order-preserving. We only have to consider three cases:
  • (1)
    $(w,t),(w^{\prime},t^{\prime})\in\mathbb{Z}_{\leq r_{1}}^{n}\times I_{q_{1}}-\mathbb{Z}_{\leq r_{1}}^{n}\times 0-\mathbb{Z}_{\leq r_{1}}^{n}\times q_{1}$ and $(w,t)\leq(w^{\prime},t^{\prime})$.
  • (2)
    $(w,q_{1}-1)\in\mathbb{Z}_{\leq r_{1}}^{n}\times(q_{1}-1)$, $(w^{\prime},q_{1})\in\mathbb{Z}_{\leq r_{1}}^{n}\times q_{1}$ and $(w^{\prime},q_{1})\leq(w,q_{1}-1)$.
  • (3)
    $(w,0)\in\mathbb{Z}_{\leq r_{1}}^{n}\times 0$, $(w^{\prime},1)\in\mathbb{Z}_{\leq r_{1}}^{n}\times 1$ and $(w,0)\leq(w^{\prime},1)$.
For case (1), $(\lambda_{r_{1}}(w),t)\leq(\lambda_{r_{1}}(w^{\prime}),t^{\prime})$ indicates that
$\lambda_{r_{1}}(|\mathcal{K}(U^{\gamma_{r_{1}}}_{w})|)\times\beta_{q_{1}}(|\mathcal{K}(U^{I_{q_{1}}}_{t})|)\subseteq\lambda_{r_{1}}(|\mathcal{K}(U^{\gamma_{r_{1}}}_{w^{\prime}})|)\times\beta_{q_{1}}(|\mathcal{K}(U^{I_{q_{1}}}_{t^{\prime}})|),$
then by the definition of $H$, $H(\lambda_{r_{1}}(w),t)\leq H(\lambda_{r_{1}}(w^{\prime}),t^{\prime})$. For case (2), since $x_{0}$ is a minimum vertex, $x_{0}\leq H(w^{\prime},q_{1})\leq H(w,q_{1}-1)$. For case (3), we have
$H(w,0)$ $=\widehat{\mu_{X}}\circ(|\mathcal{K}(f)|\circ\lambda_{r}^{-1})\circ\lambda_{r_{1}}(w)=\widehat{\mu_{X}}\circ G\circ(\lambda_{r_{1}}(w),0)$
$\subseteq\widehat{\mu_{X}}\circ G\circ\lambda_{r_{1}}(|\mathcal{K}(U^{\gamma_{r_{1}}}_{w^{\prime}})|)\times\beta_{q_{1}}(|\mathcal{K}(U^{I_{q_{1}}}_{1})|)\subseteq U_{u}^{X},$
where $H(w^{\prime},1)=u$. This says that $H(w,0)\leq H(w^{\prime},1)$.
We claim that
(3.4) $[H_{0}]=[f]$
in $\pi_{n}^{D}(X,x_{0})$. There is a sequence of partitions
$P_{r_{1}}=A_{k+1}\subseteq A_{k}\subseteq\cdots\subseteq A_{1}\subseteq A_{0}=P_{r+1},$
the former one being a refinement of the latter one by adding an $(n-1)$-dimensional hyperplane. For each $i$, construct an order-preserving map $F_{i}:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\rightarrow(X,x_{0})$ by
$F_{i}(w)=\begin{cases}\widehat{\mu_{X}}\circ(|\mathcal{K}(f)|\circ\lambda_{r}^{-1})\circ\lambda_{A_{i}}(w)~{}&\text{if}~{}w\in\gamma_{A_{i}},\\ x_{0}~{}&\text{otherwise},\end{cases}$
where $11\leq i\leq k+$ 1. If we can show that
(3.5) $[F_{i}]=[F_{i-1}],$
then this will imply
$[H_{0}]=[F_{k+1}]=\cdots=[F_{0}]=[f],$
which the claim follows. Without loss of generality we may assume the $(n-1)$-dimensional hyperplane added to $A_{i}$ from $A_{i-1}$ is $x_{1}=\frac{1}{r_{1}}$. Let $F^{j}:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\rightarrow(X,x_{0})$ be defined by
$(x_{1},x_{2},\cdots,x_{n})\mapsto\begin{cases}F_{i-1}(j,x_{2},\cdots,x_{n})~{}&\text{if}~{}x_{1}=j~{}\text{or}~{}j+1,\\ F_{i-1}(x_{1},x_{2},\cdots,x_{n})~{}&\text{if}~{}x_{1}<j,\\ F_{i-1}(x_{1}-1,x_{2},\cdots,x_{n})~{}&\text{if}~{}x_{1}>j+1,\end{cases}$
where $11\leq j\leq r+$ 1. Then
$F_{i}=F^{1}>F^{2}<F^{3}>\cdots>(<)F^{r+1}=F_{i-1}.$
Thus $[f]=[H_{0}]=0$ in $\pi_{n}^{D}(X,x_{0})$, finishing the proof of the injectivity of $\theta$.
Let $f\in\pi_{n}(|\mathcal{K}(X)|,x_{0})$. The key observation now is that under the homeomorphism $\lambda_{r}$, $\mathcal{K}(\gamma_{r})$ becomes a triangulation of $L^{n}$. To check surjectivity of $\theta$, represent $f$ by a simplicial map $f^{\prime}$ from $\mathcal{K}(\gamma_{r})$ to $\mathcal{K}(X)$ for some $r>0$, using the proof of simplicial approximation theorem [10, Lemma 14.1, Theorem 16.1]. As above, we can get an order-preserving map $g:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\rightarrow(X,x_{0})$ given by
$g(w)=\begin{cases}\widehat{\mu_{X}}\circ(|f^{\prime}|\circ\lambda_{r}^{-1})\circ\lambda_{r_{1}}(w)~{}&\text{if}~{}w\in\mathbb{Z}_{\leq r_{1}}^{n},\\ x_{0}~{}&\text{otherwise},\end{cases}$
with $r_{1}$ a multiple of $r$. $\theta(g)$ is homotopic to $|\mathcal{K}(f^{\prime})|$, because both of them agree on the boundary of each cube of $\lambda_{r}(|\mathcal{K}(\gamma_{r})|)$. ∎
We now directly compute the $1$th discrete homotopy group of a circle in the next example, leading to a new viewpoint toward the fundamental group of a circle.
Example 3.4.
Let $X$ be a poset described below with a fixed basepoint $b$, which is a finite model of the $1$-dimensional sphere $S^{1}$.
${c}$${d}$${a\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${b\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
Let $f\in\pi^{D}_{1}(X,b)$ be a map with radius $r$. Then $f$ can be represented by
$\cdots b\rightarrow x_{-r+1}\leftarrow\cdots x_{-1}\leftarrow x_{0}\rightarrow x_{1}\cdots\rightarrow x_{r-1}\leftarrow b\cdots,$
where $x_{i}$ is the value of $f$ at $i$. If two consecutive points $x_{i}$ and $x_{i+1}$ are equal, $-r+1\leq i\leq r-2$, for example, $x_{i}=x_{i+1}$ for $20\leq i\leq r-$ 2 and $i$ is even, then $f\geq f_{1}$, where $f_{1}\in\pi^{D}_{1}(X,b)$ is defined by
$\cdots b\rightarrow\cdots\leftarrow x_{0}\cdots x_{i-1}\leftarrow x_{i}\rightarrow x_{i+2}\leftarrow x_{i+2}\rightarrow x_{i+3}\cdots x_{r-1}\leftarrow b\cdots.$
Repetition of this process eventually produces a map $g\in\pi^{D}_{1}(X,b)$ given by
(3.6) $\cdots b\rightarrow y_{-n+1}\leftarrow\cdots y_{-1}\leftarrow y_{0}\rightarrow y_{1}\cdots\rightarrow y_{m-1}\leftarrow b\cdots,$
where $m,n<r$, $y_{-n+1},~{}y_{m-1}\not=b$, and two consecutive points $y_{i}$ and $y_{i+1}$, $-n\leq i\leq m-2$, are not equal.
Denote by $C_{b}$ the constant map $\mathbb{Z}\rightarrow b$. If $f(i)\not=a$ for all $i$, then
$C_{b}\leq f,$
which means $[f]=0\in\pi^{D}_{1}(X,b)$. We now assume the values of $f$, thus $g$, contain $a$. If $y_{i}=y_{i+2}=a$, then
$g$ $=\cdots a\rightarrow y_{i+1}\leftarrow a\rightarrow y_{i+3}\cdots$
$>g_{1}$ $=\cdots a\rightarrow a\leftarrow a\rightarrow y_{i+3}\cdots$
$<g_{2}$ $=\cdots a\rightarrow y_{i+3}\leftarrow y_{i+3}\rightarrow y_{i+3}\cdots$
$>g_{3}$ $=\cdots a\rightarrow y_{i+3}\leftarrow b\rightarrow y_{i+3}\cdots.$
This argument does apply for the case that $y_{i}=y_{i+2}=b$. After repeating this process, we obtain a map $h\in\pi^{D}_{1}(X,b)$ given by
(3.7) $\cdots b\rightarrow z_{-n+1}\leftarrow\cdots z_{-1}\leftarrow z_{0}\rightarrow z_{1}\cdots\rightarrow z_{m-1}\leftarrow b\cdots,$
where if $z_{i}=a$ for $-n+2\leq i\leq m-2$, then $z_{i-2}=z_{i+2}=b$, and if $z_{i}=b$ for $-n+4\leq i\leq m-4$, then $z_{i-2}=z_{i+2}$, and two consecutive vertices $z_{i}$ and $z_{i+1}$, $-n+1\leq i\leq m-2$, are distinct.
Note that $b$ appears in (3.7) with a period of four in the form
(3.8) $b\rightarrow z_{i}\leftarrow a\rightarrow z_{i+2}\leftarrow b,$
where $-n+1\leq i\leq m-3$. If $z_{i}$ and $z_{i-1}$ are equal, for example, $z_{i}=z_{i+2}=c$, then
$h$ $=\cdots b\rightarrow c\leftarrow a\rightarrow c\leftarrow b\cdots$
$<h_{1}$ $=\cdots b\rightarrow c\leftarrow c\rightarrow c\leftarrow b\cdots$
$>h_{2}$ $=\cdots b\rightarrow b\leftarrow b\rightarrow b\leftarrow b\cdots.$
This process, together with the methods of obtaining (3.6), can be repeated to get a map $k\in\pi^{D}_{1}(X,b)$ defined by
(3.9) $\cdots b\rightarrow v_{-n+1}\leftarrow\cdots v_{-1}\leftarrow v_{0}\rightarrow v_{1}\cdots\rightarrow v_{m-1}\leftarrow b\cdots,$
in which (3.8) becomes one of the two forms:
$b\rightarrow c\leftarrow a\rightarrow d\leftarrow b,$
$b\rightarrow d\leftarrow a\rightarrow c\leftarrow b,$
the numbers of them in (3.9) being denoted as $t_{1}$ and $t_{2}$, respectively.
Let $e\in\pi^{D}_{1}(X,b)$ be a map of radius $2$ given by
$\cdots b\rightarrow d\leftarrow a\rightarrow c\leftarrow b\cdots.$
Observe that
$e\cdot e^{-1}=$ $\cdots b\rightarrow d\leftarrow a\rightarrow c\leftarrow b\rightarrow c\leftarrow a\rightarrow d\leftarrow b\cdots$
$<$ $\cdots b\rightarrow d\leftarrow a\rightarrow c\leftarrow c\rightarrow c\leftarrow a\rightarrow d\leftarrow b\cdots$
$>$ $\cdots b\rightarrow d\leftarrow a\rightarrow a\leftarrow a\rightarrow a\leftarrow a\rightarrow d\leftarrow b\cdots$
$<$ $\cdots b\rightarrow d\leftarrow d\rightarrow d\leftarrow d\rightarrow d\leftarrow a\rightarrow d\leftarrow b\cdots$
$>$ $\cdots b\rightarrow b\leftarrow b\rightarrow b\leftarrow b\rightarrow b\leftarrow b\rightarrow b\leftarrow b\cdots=C_{b}.$
Similar argument implies $e^{-1}\cdot e=C_{b}$. We have just seen that $[f]=[k]$ is the product of finitely many $e$ and $e^{-1}$. With the techniques of getting (3.6), $k$ is homotopic to a map $l\in\pi^{D}_{1}(X,b)$, which is an element of the free group $F([e])$ generated by $[e]$. This says $\pi^{D}_{1}(X,b)=F([e])=\mathbb{Z}$. In particular,
$[f]=[k]=(t_{2}-t_{1})[e].$
4. Discrete homology of finite posets
4.1. Discrete cubical homology
Given a finite poset $M$. Let $Q_{1}$ denote the finite poset with vertex set $\{0,1\}$ and an edge $(0,1)$. For each $n\geq 1$, let $Q_{n}$ be the finite poset given by
$Q_{n}=\begin{cases}0~{}&\text{if}~{}n=0,\\ Q_{1}^{n}~{}&\text{if}~{}n\geq 1.\end{cases}$
Definition 4.1.
An $n$-cube of $M$ is a map $f:Q_{n}\rightarrow M$.
Let $L_{n}(M)$ be the free abelian group generated by all $n$-cubes. We define two face maps from $L_{n}(M)$ to $L_{n-1}(M)$. Let $\sigma\in L_{n}(M)$. For $n\geq 1$, $i\in\{1,\cdots,n\}$ and $(a_{1},\cdots,a_{n-1})\in Q_{n-1}$, let $f_{i}^{-}\sigma$ and $f_{i}^{+}\sigma$ be the $(n-1)$-cubes of $M$ given by
$f^{-}_{i}\sigma(a_{1},\cdots,a_{n-1})=\sigma(a_{1},\cdots,a_{i-1},0,a_{i},\cdots,a_{n-1})$
$f^{+}_{i}\sigma(a_{1},\cdots,a_{n-1})=\sigma(a_{1},\cdots,a_{i-1},1,a_{i},\cdots,a_{n-1}).$
An $n$-cube $\sigma$ is degenerate if $f_{i}^{-}=f_{i}^{+}$ for some $i$. Let $D_{n}(M)$ be the free abelian group generated by all degenerate $n$-cubes and $C_{n}(M)=L_{n}(M)/D_{n}(M)$, whose elements are called $n$-chains in $M$.
Definition 4.2.
The boundary of an $n$-cube $\sigma:Q_{n}\rightarrow M$ is
$\partial_{n}^{\rm Cube}(\sigma)=\sum_{i=1}^{n}(-1)^{i}(f_{i}^{-}\sigma-f_{i}^{+}\sigma).$
The boundary operator extends to a group homomorphism $\partial_{n}^{\rm Cube}:L_{n}(M)\rightarrow L_{n-1}(M)$. Since $\partial_{n}^{\rm Cube}(D_{n}(M))\subseteq D_{n-1}(M)$, we obtain a map $\partial_{n}^{\rm Cube}:C_{n}(M)\rightarrow C_{n-1}(M)$ and a chain complex $(C_{*}(M),\partial_{*}^{\rm Cube})$. Let
$H^{\rm Cube}_{n}(M)=\rm{Ker}\partial_{n}^{\rm Cube}/\rm{Im}\partial_{n+1}^{\rm Cube}.$
The group $H_{n}^{\rm Cube}(M)$ is called the $n$th discrete cubical homology group of $M$. If $z\in\rm{Ker}\partial_{n}^{\rm Cube}$, then we write $[z]$ for the equivalence class of $z$ in $H_{n}^{\rm Cube}(M)$.
Theorem 4.3.
Let $M_{1}$ and $M_{2}$ be finite posets. If $M_{1}$ and $M_{2}$ are homotopy equivalent, then
$H^{\rm Cube}_{n}(M_{1})=H^{\rm Cube}_{n}(M_{2})$
for all $n\geq 0$.
The idea of the proof of this theorem is exacly the same as the one in [3, Theorem 3.4]. We will not prove this theorem here.
We now construct a natural homomorphism from the discrete to the simplicial homology groups.
Let $K$ be a simplicial complex. For $n\geq 0$, we denote by $L^{\rm Simpl}_{n}(K)$ the free abelian group generated by all nondegenerate simplices of $K$ and denote by $D^{\rm Simpl}_{n}(K)$ the free abelian group generated by all degenerate simplices of $K$. Let $C^{\rm Simpl}_{n}(K)=L^{\rm Simpl}_{n}(K)/D^{\rm Simpl}_{n}(K)$. The $n$th simplicial homology group of $K$ is denoted $H_{n}^{\rm Simpl}(K)$ whose associated chain complex is $(C^{\rm Simpl}_{*}(K),\partial_{*}^{\rm Simpl})$. Similarly, $H_{*}^{\rm Sing}$ denotes the singular homology theory. Denote by $\eta_{*}$ the canonical homomorphism
$H_{n}^{\rm Simpl}(K)\rightarrow H_{*}^{\rm Sing}(|K|).$
To establish a natural homomorphism between the homology groups ${H}_{*}^{\rm Cube}(M)$ and ${H}_{*}^{\rm Simpl}(\mathcal{K}(M))$, we are motivated by the construction of the one in [3], which is a natural map from the cubical homology groups to the path homology groups. This map will be defined in a similar fashion.
Consider an $n$-cube $\sigma:Q_{n}\rightarrow M$ with $n\geq 1$. In order to define a map from ${C}_{n}^{\rm Cube}(M)$ to ${C}^{\rm Simpl}_{n}(\mathcal{K}(M))$, we associate to any permutation $\tau\in S_{n}$ an $n$-chain $p_{\tau}=\{x_{0},x_{1},\cdots,x_{n}~{}|~{}x_{0}<x_{1}<\cdots<x_{n}\}$ from $(0,\cdots,0)\in V(Q_{n})$ to $(1,\cdots,1)\in V(Q_{n})$. The chain $p_{\tau}$ is defined as follows. Let $x_{0}=(0,\cdots,0)$. For $0<i\leq n$, flipping the $\tau(i)$-th coordinate of $x_{i-1}$ from 00 to $1$ gives $x_{i}$. We write $p_{\tau}(i)$ for the $i$-th vertex in the chain $p_{\tau}$ for $0\leq i\leq n$.
To each $n$-cube $\sigma\in{L}_{n}^{\rm Cube}(M)$ for $n\geq 1$, we assign the chain
$\psi(\sigma)$ $=\sum_{\tau\in S_{n}}{\rm sign}(\tau)\mathcal{K}(\{\sigma\circ p_{\tau}(i)~{}|~{}0\leq i\leq n\})$
$=\sum_{\tau\in S_{n}}{\rm sign}(\tau)[\sigma\circ p_{\tau}(0),\sigma\circ p_{\tau}(1),\cdots,\sigma\circ p_{\tau}(n)]$
of ${L}^{\rm Simpl}_{n}(\mathcal{K}(M))$. It is not hard to see that if $\sigma\in{D}_{n}^{\rm Cube}(M)$, then $\psi(\sigma)\in{D}^{\rm Simpl}_{n}(\mathcal{K}(M))$.
The proof that $\psi$ is a chain map is similar to the one in [3, Lemma 5.1]. For completeness, we will sketch this proof.
Proposition 4.4.
Let $\sigma\in{C}_{n}^{\rm Cube}(M)$. Then
$\partial_{n}^{\rm Simpl}\psi(\sigma)=\psi(\partial_{n}^{\rm Cube}(\sigma)).$
Hence $\psi$ is a chain map and induces a homomorphism
$\psi_{*}:{H}_{p}^{\rm Cube}(M)\rightarrow{H}_{p}^{\rm Simpl}(\mathcal{K}(M))$
for each $p\geq 0$.
Proof.
If $n=0$ or $n=1$, the proof is trivial and then in what follows we assume $n\geq 2$. By the boundary formula,
(4.1) $\partial_{n}^{\rm Simpl}([\sigma\circ p_{\tau}(0),\cdots,\sigma\circ p_{\tau}(n)])=\sum_{i=0}^{n}(-1)^{i}[\sigma(p_{\tau}(0)),\cdots,\widehat{\sigma(p_{\tau}(i))},\cdots,\sigma(p_{\tau}(n))].$
Let $11\leq l\leq n-$ 1 . If $\tau^{\prime}$ is constructed from $\tau$ by interchanging $\tau(l)$ and $\tau(l+1)$, then $p_{\tau}(i)=p_{\tau^{\prime}}(i)$ for $i\not=l$. Thus the $l$-th summands of (4.1) for $p_{\tau}$ and $p_{\tau^{\prime}}$ coincide. In addition, ${\rm sign}(\tau)={\rm-sign}(\tau^{\prime})$. This means that
$\partial_{n}^{\rm Simpl}\psi(\sigma)=$ $\sum\limits_{\tau\in S_{n}}{\rm sign}(\tau)\partial_{n}^{\rm Simpl}([\sigma\circ p_{\tau}(0),\cdots,\sigma\circ p_{\tau}(n)])$
$=$ $\sum\limits_{\tau\in S_{n}}{\rm sign}(\tau)([\sigma(p_{\tau}(1)),\cdots,\sigma(p_{\tau}(n))]$
$+(-1)^{n}[\sigma(p_{\tau}(0)),\cdots,\sigma(p_{\tau}(n-1))]).$
To each $\tau\in S_{n}$, we associate the following two permutaions.
  • (1)
    $\tau^{-}\in S_{n-1}$ is defined as the permutation where $\tau^{-}(j)=\tau(j)$ if $\tau(j)<\tau(n)$ and $\tau(j)-1$ if $\tau(j)>\tau(n)$.
  • (2)
    $\tau^{+}\in S_{n-1}$ is defined as the permutation where $\tau^{+}(j)=\tau(j+1)$ if $\tau(j+1)<\tau(1)$ and $\tau(j+1)-1$ if $\tau(j+1)>\tau(1)$.
Then we have
$\psi(f_{i}^{-}\sigma)$ $=\sum\limits_{\tau\in S_{n},\tau(n)=i}{\rm sign}(\tau^{-})[(f_{i}^{-}\sigma)\circ(p_{\tau^{-}}(0)),\cdots,(f_{i}^{-}\sigma)\circ(p_{\tau^{-}}(n))]$
$=(-1)^{n-i}\sum\limits_{\tau\in S_{n},\tau(n)=i}{\rm sign}(\tau)[(f_{i}^{-}\sigma)\circ(p_{\tau^{-}}(0)),\cdots,(f_{i}^{-}\sigma)\circ(p_{\tau^{-}}(n))],$
$\psi(f_{i}^{+}\sigma)$ $=\sum\limits_{\tau\in S_{n},\tau(1)=i}{\rm sign}(\tau^{+})[(f_{i}^{+}\sigma)\circ(p_{\tau^{+}}(0)),\cdots,(f_{i}^{+}\sigma)\circ(p_{\tau^{+}}(n))]$
$=(-1)^{i-1}\sum\limits_{\tau\in S_{n},\tau(1)=i}{\rm sign}(\tau)[(f_{i}^{+}\sigma)\circ(p_{\tau^{+}}(0)),\cdots,(f_{i}^{+}\sigma)\circ(p_{\tau^{+}}(n))].$
Finally, observe that for $j=\tau(n)$ and $j^{\prime}=\tau(1)$,
(4.2) $\begin{split}\sigma(p_{\tau}(m))&=(f_{j}^{-}\sigma)\circ(p_{\tau^{-}}(m)),\\ \sigma(p_{\tau}(r))&=(f_{j^{\prime}}^{+}\sigma)\circ(p_{\tau^{+}}(r-1)),\end{split}$
where $0\leq m\leq n-1,~{}1\leq r\leq n$. Thus we get
[…]
These two homology theories will turn out to be closely related by the homomorphism $\psi_{*}$ in what follows.
Proposition 4.5.
Let $M$ be a homogeneous poset of dimension $1$. The map $\psi$ induces an isomorphism $\psi_{*}:{H}_{p}^{\rm Cube}(M)\rightarrow{H}_{p}^{\rm Simpl}(\mathcal{K}(M))$ for any $p\geq 0$.
Proof.
The case $p=0$ is trivial. The dimension of every maximal chain of $M$ is not greater than $1$, then this says $C_{p}^{\rm Simpl}(M)=0$ for $p>1$, so ${H}_{p}^{\rm Simpl}(\mathcal{K}(M))=0$ for $p>1$. It is not difficult to see that every $p$-cube with $p>1$ is degenerate, which means $C_{p}^{\rm Cube}(M)=0$. Then ${H}_{p}^{\rm Cube}(\mathcal{K}(M))=0$ for $p>1$.
Consider $p=1$. A nondegenerate $1$-cube can be represented by a directed edge $(v_{1},v_{2})$ of the Hasse diagram of $M$, while a nondegenerate $1$-simplex is $\{v_{1},v_{2}\}$ where $v_{1}$ and $v_{2}$ are vertices of $M$. Hence $C_{1}^{\rm Cube}(M)$ has the same basis as $C_{1}^{\rm Simpl}(M)$. Because the differentials $\partial_{1}^{\rm Cube}$ and $\partial_{1}^{\rm Simpl}$ are indeed identical, $\psi_{*}$ is an isomorphism for $p=1$. ∎
When $M$ is a poset with $|\mathcal{K}(M)|$ being an $m$-manifold, Proposition 4.5 would generalize to higher dimensions. Before proceeding further we need a useful excision property for finite posets.
Lemma 4.6.
Let $M$ be a finite poset and let $L$ be a subposet of $M$. Then $|\mathcal{K}(M-L)|$ is a strong deformation retract of $|\mathcal{K}(M)|-|{\mathcal{K}(L)}|$.
The argument given in [1, Theorem 1.4.6] works just as well for this case.
Theorem 4.7.
Let $M$ be a poset with $|\mathcal{K}(M)|$ being a closed $m$-manifold. Then the homomorphism $\psi_{*}:{H}_{p}^{\rm Cube}(M)\rightarrow{H}_{p}^{\rm Simpl}(\mathcal{K}(M))$ is a surjection for $p\not=m$.
Proof.
We prove the theorem by induction on the number $n$ of generators of chain complex $\mathcal{C}_{*}^{\rm Cube}(M)$. If $n=1$, then $M$ consists of a single vertex $v$. In each dimension $p$, there is exactly one $p$-cube $g:Q_{p}\rightarrow v$ of $M$ and exactly one ordered $p$-simplex $\{v,\cdots,v\}$ of $\mathcal{K}(M)$. Furthermore, $\psi(g)=\{v,\cdots,v\}$. Then $\psi$ is an isomorphism already on the chain level. Hence $\psi_{*}$ is surjective in this case.
Now suppose the theorem holds for any finite poset whose associated cubical chain complex has fewer than $n$ generators. Let $M$ be a poset with ${C}_{*}^{\rm Cube}(M)$ having $n$ generators. Given $p\geq 0$. If $p>m$, then
${H}_{p}^{\rm Simpl}(\mathcal{K}(M))=0,$
where the equality follows from [6, Lemma 3.27]. Therefore $\psi_{*}$ is surjective for $p>m$.
The case $p=0$ is trivial. Assume that $10<p\leq m-$ 1. Let $g:Q_{p}\rightarrow M$ be a generator of ${C}_{p}^{\rm Cube}(M)$. If $g(Q_{p})=M$, then
$H_{p}^{\rm Simpl}(\mathcal{K}(M))=H_{p}^{\rm Simpl}(\mathcal{K}(g(Q_{p})))=0$
since $g(Q_{p})$ is contractible, so we are done. Now assume $g(Q_{p})\not=M$. Consider the following commutative diagram, where $i_{*},j_{*}$ are induced by inclusion maps.
[…]
Note that $|\mathcal{K}(M-g(Q_{p}))|$ is an $m$-manifold, being a strong deformation retract of $|\mathcal{K}(M)|-|{\mathcal{K}(g(Q_{p}))}|$ by Lemma 4.6. It follows from the induction hypothesis that the map $\psi_{*}$ on the bottom line of the diagram is surjective for $p\leq m-1$. We shall show that $j_{*}$ is a surjection for $p\leq m-1$, which means that the map $\psi_{*}$ on the top line has the same property and the proof is complete.
Because $g(Q_{p})$ is contractible, ${\mathcal{K}(g(Q_{p}))}$ is collapsible [1, Theorem 4.2.11]. Let us take $N$ to be a regular neighbourhood of $|{\mathcal{K}(g(Q_{p}))}|$ in $|\mathcal{K}(M)|$. It is known that $N$ is a PL $m$-ball [11, Corollary 3.27]. By excision, we have
${H}_{*}^{\rm Sing}(|\mathcal{K}(M)|,|\mathcal{K}(M)|-|{\mathcal{K}(g(Q_{p}))}|)={H}_{*}^{\rm Sing}({\rm int}N,{\rm int}N-|{\mathcal{K}(g(Q_{p}))}|).$
Let $h:{\rm int}N\rightarrow\mathbb{R}^{m}$ be a homeomorphism. Then
${H}_{*}^{\rm Sing}({\rm int}N,{\rm int}N-|{\mathcal{K}(g(Q_{p}))}|)$ $={H}_{*}^{\rm Sing}((\mathbb{R}^{m},\mathbb{R}^{m}-h(|{\mathcal{K}(g(Q_{p}))}|)),$
${H}_{*}^{\rm Sing}({\rm int}N-|{\mathcal{K}(g(Q_{p}))}|)$ $={H}_{*}^{\rm Sing}(\mathbb{R}^{m}-h(|{\mathcal{K}(g(Q_{p}))}|)).$
If we take $B^{p}\subseteq\mathbb{R}^{m}$ to be a standard $p$-ball with radius sufficiently small, then Alexander duality gives
${H}_{*}^{\rm Sing}(\mathbb{R}^{m}-h(|{\mathcal{K}(g(Q_{p}))}|)={H}_{*}^{\rm Sing}(\mathbb{R}^{m}-B^{p}).$
It follows that ${H}_{i}^{\rm Sing}(|\mathcal{K}(M)|,|\mathcal{K}(M-g(Q_{p}))|)=\mathbb{Z}$ for $i=m$ and vanishes otherwise. Hence
$H_{p}^{\rm Sing}(|\mathcal{K}(M-g(Q_{p}))|)=H_{p}^{\rm Sing}(|\mathcal{K}(M)|-|{\mathcal{K}(g(Q_{p}))}|).$
The long exact sequence of homology groups for the pair $(|\mathcal{K}(M)|,|\mathcal{K}(M-g(Q_{p}))|)$ then implies $j_{*}$ has the desired property. ∎
In fact, we get a stronger result in this proof.
Corollary 4.8.
Let $p\leq m-1$. Then $\psi_{*}:{H}_{p}^{\rm Cube}(M-g(Q_{p}))\rightarrow{H}_{p}^{\rm Simpl}(\mathcal{K}(M))$ is a surjection for any $g\in L_{p}(M)$ with $|\mathcal{K}(g(Q_{p}))|\subseteq{\rm int}|\mathcal{K}(M)|$, where $M$ is a poset with $|\mathcal{K}(M)|$ being an $m$-manifold.
Theorem 4.7 might suggest conjecturing that $H_{*}^{\rm Cube}(M)$ and $H_{*}^{\rm Simpl}(\mathcal{K}(M))$ are the same for all finite posets $M$. Now we construct an example showing this is not always the case.
Example 4.9.
Let $M$ be the finite poset described below.
${e}$${f}$${c\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${d\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${a\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${b\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
This is a finite model of the $2$-dimensional sphere, hence
$H_{1}^{\rm Simpl}(\mathcal{K}(M))=0.$
Represent $1$-cubes $f:Q_{1}\rightarrow M$ by $(x_{0},x_{1})$, where $x_{i}=f(i)$, and represent $2$-cubes $f:Q_{2}\rightarrow M$ by $(x_{00},x_{10},x_{01},x_{11})$, where $x_{ij}=f(i,j)$. We shall prove $H_{1}^{\rm Cube}(M)\not=0$ by finding a $1$-cycle $\alpha\in C^{\rm Cube}_{1}(M)$ that is not a boundary. Define
$\alpha=(a,e)-(b,e)+(b,f)-(a,f).$
It is easy to check $\alpha$ is a cycle.
If $\beta\in C_{2}^{\rm Cube}(M)$ satisfies $\partial_{2}^{\rm Cube}(\beta)\not=0$, then $\beta$ is a linear combination
$x_{1}(a,c,a,e)+x_{2}(a,d,a,e)+x_{3}(a,d,c,e)$
$+$ $x_{4}(a,c,a,f)+x_{5}(a,d,a,f)+x_{6}(a,d,c,f)$
$+$ $x_{7}(b,c,b,e)+x_{8}(b,d,b,e)+x_{9}(b,d,c,e)$
$+$ $x_{10}(b,c,b,f)+x_{11}(b,d,b,f)+x_{12}(b,d,c,f).$
If $\alpha$ is a boundary, there is a $\beta\in C_{2}^{\rm Cube}(M)$ such that $\partial_{2}^{\rm Cube}(\beta)=\alpha.$ Then we have
$\begin{cases}x_{1}-x_{3}+x_{7}-x_{9}=0\\ x_{1}-x_{3}+x_{4}-x_{6}=0\\ -x_{1}-x_{2}=1\\ x_{2}+x_{3}+x_{8}+x_{9}=0\\ x_{2}+x_{3}+x_{5}+x_{6}=0\\ x_{4}-x_{6}+x_{10}-x_{12}=0\\ -x_{4}-x_{5}=-1\\ x_{5}+x_{6}+x_{11}+x_{12}=0\\ x_{7}-x_{9}+x_{10}-x_{12}=0\\ -x_{7}-x_{8}=-1\\ x_{8}+x_{9}+x_{11}+x_{12}=0\\ -x_{10}-x_{11}=1.\end{cases}$
Observe that
$(-x_{10}-x_{11})=$
$-(\left(x_{7}-x_{9}+x_{10}-x_{12})+(-x_{7}-x_{8})+(x_{8}+x_{9}+x_{11}+x_{12})\right).$
It follows that $\beta=0$, a contradiction.
4.2. An application of Corollary 4.8
As we have seen, for a given poset $M$, some information about the homology groups of $M$ can be squeezed out of Corollary 4.8 by computing discrete cubical homology groups of a subposet $M$ with fewer vertices. In this subsection, going a step further, we generalize Corollary 4.8 in the context of a class of finite posets.
Let $M$ be a homogeneous poset of dimension $m$. Define $M^{=p}$ as the subposet of elements of degree equal to $p$,
$M^{=p}=\{m^{p}_{1},\cdots,m^{p}_{i_{p}}|~{}m^{p}_{i_{j}}\in M,{\rm deg}(m^{p}_{i_{j}})=p\}.$
Label the vertices of $M$ by
(4.3) $M=\bigcup_{\begin{subarray}{c}1\leq t\leq i_{p}\\ 0\leq p\leq m\end{subarray}}m^{p}_{t}.$
It is no difficult to see that if $M$ is a poset with $|\mathcal{K}(M)|$ being an $m$-manifold, then $M$ is homogeneous.
Let $A=\mathcal{K}(\{m^{0}_{t_{0}},\cdots,m^{m}_{t_{m}}\})$ be an $m$-simplex of $\mathcal{K}(M)$ and let $I^{m}=[0,1]^{m}$ be the $m$-cube in $\mathbb{R}^{m}$. The last coordinate of $I^{m}$ is singled out for special reference and written as $I^{m}=I^{m-1}\times I$. Regard $|A|$ as a polyhedron of $I^{m}$ by the canonical homeomorphism $f$ given by
$f(m_{t_{i}})=\begin{cases}(0,0\cdots,0),&~{}\text{if}~{}i=0,\\ e_{i},&~{}\text{if}~{}1\leq i\leq m-1,\\ (\frac{1}{m},\frac{1}{m}\cdots,\frac{1}{m}),&~{}\text{if}~{}i=m.\end{cases}$
If $(x,t),~{}(x^{\prime},t^{\prime})\in|A|\subseteq I^{m-1}\times I$, we say $(x,t)$ is directly below $(x^{\prime},t^{\prime})$ in $|A|$ if $x=x^{\prime}$ and $t<t^{\prime}$.
Definition 4.10.
Let $a,~{}a^{\prime}\in|\mathcal{K}(M)|$. We say $a$ is in the shadow of $a^{\prime}$ in $|\mathcal{K}(M)|$ if there is an $m$-chain $A$ of $M$ such that $|\mathcal{K}(A)|$ contains both $a$ and $a^{\prime}$, and $a$ is directly below $a^{\prime}$ in $|\mathcal{K}(A)|$.
Definition 4.11.
Let $X$ be a subpolyhedron of $|\mathcal{K}(M)|$. The shadow ${\rm sh}(X)$ of $X$ is the closure of the set of points of $X$ that are directly below some other point of $X$.
Given polyhedra $X\subseteq Y\subseteq|\mathcal{K}(M)|$ such that $Y\searrow X$ is an elementary collapse, we call this collapse sunny if no points of $Y-X$ are in ${\rm sh}(Y)$. If $X$ is a point, then $Y$ is called sunny collapsible.
Lemma 4.12.
Let $M$ be a poset with $|\mathcal{K}(M)|$ being an $m$-manifold and $f\in L_{n}(M)$. If ${\dim}f(Q_{n})<m$, then $|\mathcal{K}(f(Q_{n}))|$ is suuny collapsible.
Proof.
Denote by $v$ the image of $(0,\cdots,0)$ under $\mathcal{K}(f)$. We shall construct inductively a decreasing sequence of subpolyhedra
$|\mathcal{K}(f(Q_{n}))|=X_{0}\supseteq X_{1}\supseteq\cdots\supseteq X_{k}=v,$
satisfying the two conditions:
  • (1)
    $X_{i}=|vK^{n-i-1}|$, where $K^{n-i-1}$ is a principle $(n-i-1)$-complex.
  • (2 )
    There is a sunny collapse $X_{i-1}\searrow X_{i}$.
Let $K^{n-1}=\{A\in\mathcal{K}(f(Q_{n}))~{}|~{}v~{}\text{is not a vertex of}~{}A\}$. Then $X_{0}=|vK^{n-1}|$.
Suppose we are given $,X_{i-1},K^{n-i}$, satisfying the two conditions, we have to construct $X_{i},~{}K^{n-i-1}$ and prove the two conditions. Define
$K^{n-i-1}=\text{the}~{}(n-i-1)~{}\text{- skeleton of}~{}K^{n-i},$
which is a principle complex since $K^{n-i}$ is principle. Let $X_{i}=|vK^{n-i-1}|$.
For each $(n-i)$-simplex $D\in K^{n-i}$, the closure ${\rm cl}(v|D|)$ is a ball and $|D|$ is a face. Collapsing each ${\rm cl}(v|D|)$ from ${\rm cl}(|D|)$, we obtain
$|vK^{n-i}|\searrow|vK^{n-i-1}|,$
which is a sunny collapse because no points of ${\rm sh}(X_{i-1})$ are removed. ∎
[12, Lemma 20] states that if $(B^{m},B^{q})$ is an $(m,q)$-ball pair of codimension $m-q\geq 3$, then $B^{m}\searrow B^{q}$. If $M$ is a contractible poset with $|\mathcal{K}(M)|$ being an $m$-manifold, then $|\mathcal{K}(M)|$ is a PL ball since a collapsible PL manifold is a PL ball [11, Corollary 3.27]. Note that $(|\mathcal{K}(M)|,|\mathcal{K}(f(Q_{n}))|)$ may be not a ball pair [12], since $\partial|\mathcal{K}(f(Q_{n}))|$ and ${\rm int}|\mathcal{K}(f(Q_{n}))|$ are not necessarily contained in $\partial|\mathcal{K}(M)|$ and ${\rm int}|\mathcal{K}(M)|$, respectively. Moreover, the codimension may be less than $3$. Thus we cannot apply [12, Lemma 20] to show that $|\mathcal{K}(M)|\searrow|\mathcal{K}(f(Q_{n}))|$ if ${\dim}f(Q_{n})<m$. However, the techniques in this proof can be used to obtain our result.
Lemma 4.13.
Let $M$ be a contractible poset with $|\mathcal{K}(M)|$ being an $m$-manifold and $f\in L_{n}(M)$. Then $|\mathcal{K}(M)|\searrow|\mathcal{K}(f(Q_{n}))|$ if ${\dim}f(Q_{n})<m$.
Proof.
The sunny collapse in the proof of Lemma 4.12 is simplicial, written as
[…]
Let $L_{i}$ be the polyhedron consisting of $|\mathcal{K}(f(Q_{n}))|$ and all points below $|K_{i}|$ in $|\mathcal{K}(M)|$. We will prove that
$|\mathcal{K}(M)|\searrow L_{0}\searrow L_{1}\searrow\cdots\searrow L_{r}\searrow|K|.$
Choose a triangulation $Q$ of $|\mathcal{K}(M)|$ containing a subdivision $Q_{0}$ of $L_{0}$. The collapsibility of $|\mathcal{K}(M)|$ allows us to collapse $Q$ to $Q_{0}$ simplicially from the top, in order of decreasing dimension of simplices, completing the first step. The last step is obvious. If ${\rm deg}f(0,\cdots,0)<m$, then $L_{r}=|K|$. If ${\rm deg}f(0,\cdots,0)=m$, then $L_{r}$ consists of $K$ joined by some single arcs formed by points below $f(0,\cdots,0)$ in $|\mathcal{K}(M)|$. Collapsing these arcs finishes this step.
[…]
$m^{p}_{t}|\mathcal{K}\{m^{p_{1}}_{t_{1}},\cdots,m^{p_{k}}_{t_{k}}\}|,$
where $a=m_{t}^{p},~{}p_{1}<\dots<p_{k},~{}1\leq t_{j}\leq i_{p_{j}}$, and $|B|=|\mathcal{K}\{m^{p_{1}}_{t_{1}},\cdots,m^{p_{k}}_{t_{k}}\}|$. If
$L(|B|)=L(|A|)=\emptyset,$
we are done.
Now assume that
$L(|B|)\not=\emptyset,$
this is equivalent to saying that $p_{k}=m$. It follows that
$L(b)\not=\emptyset,$
for any $b\in{\rm int}|B|$. In particular, $L(\widehat{B})\not=\emptyset$, where $\widehat{B}$ is the barycenter of $|B|$. Let $C_{1},\cdots,C_{r}$ be all $m$-chains of $M$ containing $\{m^{p_{1}}_{t_{1}},\cdots,m^{p_{k}}_{t_{k}},m^{p}_{t}\}$. Let us take a point $c_{j}\in|\mathcal{K}(C_{j}-m^{p_{k}}_{t_{k}})|$ for each $j$ that is directly below the barycenter of $B$. If $c_{j}\in|\mathcal{K}(f(Q_{n}))|$, then
$|\mathcal{K}(C_{j}-m^{p_{k}}_{t_{k}})|\subseteq|\mathcal{K}(f(Q_{n}))|,$
since $c_{j}$ is by definition in the interior of $|\mathcal{K}(C_{j}-m^{p_{k}}_{t_{k}})|$. This forces
$m^{p_{k}}_{t_{k}}|\mathcal{K}(C_{j}-m^{p_{k}}_{t_{k}})|\subseteq|\mathcal{K}(f(Q_{n}))|,$
$m^{p_{k}}_{t_{k}}$ being a vertex of $f(Q_{n})$, which contradicts the dimension of $|\mathcal{K}(f(Q_{n}))|$. Hence $c_{j}\not\in|\mathcal{K}(f(Q_{n}))|$ and
$|c_{j}A|\cap|\mathcal{K}(f(Q_{n}))|=|A|.$
Let $T_{j}=L_{i}\cup c_{j}A$. Then
$L_{i-1}=\bigcup\limits_{j}T_{j}.$
No points of $|K_{i}|$ are above $|A|$ since this collapse is sunny. It follows that $L_{i}\cap|c_{j}A|=|m^{p}_{t}\partial(c_{j}B)|$, $L_{i}\cup|c_{j}A|=T_{j}$. Each $T_{j}\searrow L_{i}$ is obtained by collapsing $|c_{j}A|$ from $|c_{j}B|$. This finishes the proof that $L_{i-1}\searrow L_{i}$ in this case.
It remains to consider the case
$L(|B|)=\emptyset,~{}L(|A|)\not=\emptyset.$
This says $p=m,~{}p_{k}<m$. We take a point $c^{\prime}_{j}\in|\mathcal{K}(C_{j}-m_{t}^{p})|$ for each $j$ that is directly below $m_{t}^{p}$. The argument for the previous case does apply in this case with $c_{j}$ replaced by $c_{j}^{\prime}$. ∎
Figure 2 shows a poset $M$ and the geometric realization $|\mathcal{K}(M)|$. To illustrate this proof, let $f\in L_{1}(M)$ be represented by $(C,D)$. Then the shaded area in the figure on the right is $L_{0}$.
Figure 2.
Observe that an $(m,q)$-ball pair $(B^{m},B^{q})$ is unknotted by the argument in the proof of [12, Theorem 9], provided that $B^{m}\searrow B^{q}$. If $\partial|\mathcal{K}(f(Q_{n}))|\subseteq\partial|\mathcal{K}(M)|$ and ${\rm int}|\mathcal{K}(f(Q_{n}))|\subseteq{\rm int}|\mathcal{K}(M)|$, $(|\mathcal{K}(M)|,|\mathcal{K}(f(Q_{n}))|)$ is a ball pair. Thus by Lemma 4.13, the following result holds.
Proposition 4.14.
Let $M$ be a contractible poset with $|\mathcal{K}(M)|$ being an $m$-manifold and $f\in L_{n}(M)$. If ${\rm dim}f(Q_{n})<m$ and $(|\mathcal{K}(M)|,|\mathcal{K}(f(Q_{n}))|)$ is a ball pair, then
$(|\mathcal{K}(M)|,|\mathcal{K}(f(Q_{n}))|)$
is an unknotted ball pair.
We are now in a position to state and prove the main result of this subsection.
Theorem 4.15.
Let $M$ be a contractible poset with $|\mathcal{K}(M)|$ being an $m$-manifold and $f_{1}\in L_{n}(M)$ with $|\mathcal{K}(f_{1}(Q_{n}))|\subseteq{\rm int}|\mathcal{K}(M)|$. Assume that ${\rm dim}f_{1}(Q_{n})<m$ and $(|\mathcal{K}(M)|,|\mathcal{K}(f(Q_{n}))|)$ is a ball pair. Let $c=dim|\mathcal{K}(M-f_{1}(Q_{n}))|-1$. If $n\leq m-1$ and $p\leq c$, then for any $f_{2}\in L_{p}(M-f_{1}(Q_{n}))$ with $|\mathcal{K}(f_{2}(Q_{p}))|\subseteq{\rm int}|\mathcal{K}(M-f_{1}(Q_{n}))|$ and $x\in f_{1}(Q_{n})$, the $p$th Betti number of
$M-f_{1}(Q_{n})-f_{2}(Q_{p})$
for discrete cubical homology theory is nonzero if $p=m-2\leq c,~{}0<n<m-1$, or $p=m-1\leq c,~{}n=0$, or $p=0$.
Proof.
By Corollary 4.8,
$\psi_{*}:{H}_{p}^{\rm Cube}(M-f_{1}(Q_{n})-f_{2}(Q_{p}))\rightarrow{H}_{p}^{\rm Simpl}(\mathcal{K}(M-f_{1}(Q_{n})))$
is surjective. Because $(|\mathcal{K}(M)|,|\mathcal{K}(f_{1}(Q_{n}))|)$ in an unknotted ball pair,
${H}_{q}^{\rm Sing}(|\mathcal{K}(M)|-|\mathcal{K}(f_{1}(Q_{n}))|)=\begin{cases}\mathbb{Z},~{}&\text{if}~{}q=m-2,~{}0<n<m-1,~{}\text{or}\\ &~{}q=m-1,~{}n=0,~{}\text{or}~{}q=0,\\ 0,~{}&\text{otherwise}.\end{cases}$
The conclusion of this lemma is then immediate from Lemma 4.6 and Theorem 4.7. ∎
Example 4.16.
Let $M$ be the following poset, which is a finite model of the $3$-ball.
${i}$${f\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${g\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${h\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${c\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${d\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${e\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${a\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${b\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$.
Let $f_{1}\in L_{0}(M)$ be the constant map $Q_{0}\rightarrow i$, and let $f_{2}\in L_{1}(M-f_{1}(Q_{0}))$ be the represented $(b,c)$. Then $M-f_{1}(Q_{0})-f_{2}(Q_{1})$ is a finite model of the circle, being not contractible.
An interesting application of Theorem 4.15 is as follows. Let $M=|K|$ be a triangulated $m$-manifold with $K$ simplicially collapsible. Then $\chi(K)$ is collapsible [1, Theorem 4.2.11] and $|\mathcal{K}(\chi(K))|=|K|$ is an $m$-manifold. Theorem 4.7 and Theorem 4.15 tell us when the contractibility is not preserved by the manifold $M$ with two specific balls removed.
Corollary 4.17.
Let $M=|K|$ be a triangulated $m$-manifold with $K$ simplicially collapsible and $f_{1}\in L_{n}(\chi(K))$ with $|\mathcal{K}(f_{1}(Q_{n}))|\subseteq{\rm int}M$. Assume also that ${\rm dim}f_{1}(Q_{n})<m$ and $(M,|\mathcal{K}(f(Q_{n}))|)$ is a ball pair. Let $c=dim|\mathcal{K}(\chi(K)-f_{1}(Q_{n}))|-1$ If $n\leq m-1$, then for any $f_{2}\in L_{p}(\chi(K)-f_{1}(Q_{n}))$ with $|\mathcal{K}(f_{2}(Q_{p}))|\subseteq{\rm int}|\mathcal{K}(\chi(K)-f_{1}(Q_{n}))|$,
$M-|\mathcal{K}(f_{1}(Q_{n}))|-|\mathcal{K}(f_{2}(Q_{n}))|$
in not contractible, if $10<p=m-2\leq c,~{}0<n<m-$ 1, or $00<p=m-1\leq c,~{}n=$ 0.
5. Discrete Hurewicz map
The aim of this section is to construct an analogue of the Hurewicz homomorphism connecting discrete homotopy theory with discrete cubical homology theory. The discrete Hurewicz homomorphism that we define is similar to the one in [8]. As stated in Theorem 5.5, this discrete Hurewicz homomorphism indeed coincides with the classical one,
Given a finite poset $X$ with basepoint $x_{0}$. Let $f:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\rightarrow(X,x_{0})$ be a map of radius $r_{f}$. Define $H:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\times I_{r}\rightarrow(X,x_{0})$ by $H(x,i)=f(x_{1}-i,\cdots,x_{n}-i)$ for $(x,i)\in\mathbb{Z}^{n}\times I_{r}$, where $r>r_{f}$. Then $H$ is a basepoint-preserving-homotopy from $f$ to a map $H_{r}$. Take $T_{r}f:(I_{2r}^{n},\partial I_{2r}^{n})\rightarrow(X,x_{0})$ to be the restriction of $H_{r}$ to $I^{n}_{2r}$.
To each map $f:I_{p_{1}}\times\cdots\times I_{p_{n}}\rightarrow X$, we associate an element $\varphi(f)$ of $\mathcal{C}_{n}^{\rm Cube}(X)$ in the following way. If $p_{i}=0$ for some $i$, then $\varphi(f)=0$. Assume $p_{i}>0$ for all $i$. For each $x\in I_{p_{1}-1}\times\cdots\times I_{p_{n}-1}$, let $f^{x}:Q_{n}\rightarrow X$ be the $n$-cube given by
$f^{x}(y)=f(x+y),$
and then $\varphi(f)$ is defined by
$\varphi(f)=\sum_{x\in I_{p_{1}-1}\times\cdots\times I_{p_{n}-1}}f^{x}.$
Definition 5.1.
The discrete Hurewicz map $h^{D}:\pi^{D}_{n}(X,x_{0})\rightarrow{H}^{\rm Cube}_{n}(X)$ is defined by
$h^{D}([f])=[\varphi(T_{r}f)],$
where $f:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\rightarrow(X,x_{0})$ is a map of radius $r$.
This definition is justified by the following lemma.
Lemma 5.2.
  • (1)
    Assume $f:(I^{n}_{r},\partial I_{r}^{n})\rightarrow(X,x_{0})$ is a map and $\widetilde{f}:(I^{n}_{\widetilde{r}},\partial I_{\widetilde{r}}^{n})\rightarrow(X,x_{0})$ is an extension of $f$ defined by
    (5.1) $\widetilde{f}(x)=\begin{cases}f(x)~{}&\text{if}~{}x\in I^{n}_{r},\\ x_{0}~{}&\text{if}~{}x\not\in I^{n}_{r}.\end{cases}$
    Then $\varphi(f)=\varphi(\widetilde{f})$.
  • (2)
    If $f:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\rightarrow(X,x_{0})$ is a map of radius $r$, then
    $\varphi(T_{r}f)\in{\rm Ker}(\partial^{\rm Cube}_{n}).$
  • (3)
    If $f$ and $g$ are homotopic maps $(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\rightarrow(X,x_{0})$, then
    $[\varphi(T_{r_{f}}f)]=[\varphi(T_{r_{g}}g)]$
    in ${H}^{\rm Cube}_{n}(X)$, where $r_{f}$ and $r_{g}$ are radii of $f$ and $g$, respectively.
Proof.
By the definition of $\varphi$,
$\begin{split}\varphi(\widetilde{f})&=\sum_{x\in I^{n}_{r-1}}\widetilde{f}^{x}+\sum_{x\in I^{n}_{\widetilde{r}-1}-I^{n}_{r-1}}\widetilde{f}^{x}\\ &=\sum_{x\in I^{n}_{r-1}}f^{x}=\varphi(f),\end{split}$
where the second equality comes from the fact that $\widetilde{f}^{x}$ is a degenerate cube for each $I^{n}_{\widetilde{r}-1}-I^{n}_{r-1}$. This proves $(1)$.
For $1\leq i\leq n$, let $\Gamma_{i}=\{x\in I_{2r-1}^{n}~{}|~{}x_{i}=0\}$. One has
$\begin{split}\partial_{n}^{\rm Cube}\varphi(T_{r}f)&=\sum_{i=1}^{n}(-1)^{i}\sum_{x\in\Gamma_{i}}\sum_{r=0}^{2r-1}\left(f_{i}^{-}(T_{r}f)^{x+re_{i}}-f_{i}^{+}(T_{r}f)^{x+re_{i}}\right)\\ &=\sum_{i=1}^{n}(-1)^{i}\sum_{x\in\Gamma_{i}}\left(f_{i}^{-}(T_{r}f)^{x}-f_{i}^{+}(T_{r}f)^{x+(2r-1)e_{i}}\right),\end{split}$
where $e_{i}\in\mathbb{Z}^{n}$ ia a vector with $1$ in the $i$-th position and 00 everywhere else. Observe that for each $i$ and $x\in\Gamma_{i}$, $f_{i}^{-}(T_{r}f)^{x}-f_{i}^{+}(T_{r}f)^{x+(2r-1)e_{i}}$ is 00 since $T_{r}f$ is constant on $\partial I_{2r}^{n}$. This proves $(2)$.
Given a homotopy $H:\mathbb{Z}^{n}\times I_{p}\rightarrow X$ from $f$ to $g$, let $TH$ be the map given by
$\begin{split}I^{n}_{2r}\times I_{p}&\rightarrow X,\\ (x,t)&\mapsto H(x-r,t),\end{split}$
$r$ being taken the maximum integer of radii of $H_{i},~{}0\leq i\leq p$. Let $\Gamma_{i}=\{x\in I^{n}_{2r-1}\times I_{p-1}~{}|~{}x_{i}=0\}$ and let $e_{i}\in\mathbb{Z}^{n+1}$ be the vector with $1$ in the $i$-th position and 00 everywhere else. We have
$\begin{split}\partial_{n+1}^{\rm Cube}\varphi(TH)(y)&=\sum_{i=1}^{n}(-1)^{i}\sum_{x\in\Gamma_{i}}\sum_{r=0}^{2r-1}\left(f_{i}^{-}(TH)^{x+re_{i}}(y)-f_{i}^{+}(TH)^{x+re_{i}}(y)\right)\\ +&(-1)^{n+1}\sum_{x\in\Gamma_{n+1}}\sum_{r=0}^{p-1}\left(f_{n+1}^{-}(TH)^{x+re_{n+1}}(y)-f_{n+1}^{+}(TH)^{x+re_{n+1}}(y)\right)\\ &=(-1)^{n+1}\sum_{x\in\Gamma_{n+1}}\left(f_{n+1}^{-}(TH)^{x}(y)-f_{n+1}^{+}(TH)^{x+(p-1)e_{n+1}}(y)\right)\\ &=(-1)^{n+1}\sum_{x\in\Gamma_{n+1}}(TH)(x_{1}+y_{1},\cdots,x_{n}+y_{n},0)\\ -&(-1)^{n+1}\sum_{x\in\Gamma_{n+1}}(TH)(x_{1}+y_{1},\cdots,x_{n}+y_{n},p)\\ &=(-1)^{n+1}\left(\varphi({(TH)}_{0})-\varphi({(TH)}_{p})\right)(y)\\ &=(-1)^{n+1}\left(\varphi(T_{r_{f}}f)-\varphi(T_{r_{g}}g)\right)(y),\end{split}$
where the second equality follows just as in the proof of $(2)$ and the last equality follows from $(1)$. ∎
Proposition 5.3.
The discrete Hurewicz map $h^{D}:\pi^{D}_{n}(X,x_{0})\rightarrow{H}^{\rm Cube}_{n}(X)$ is a homomorphism.
Proof.
For any two maps $f,g:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\rightarrow(X,x_{0})$, $\varphi(T_{r_{f}+r_{g}}(f\cdot g))-(\varphi(T_{r_{f}}f)+\varphi(T_{r_{g}}g))$ is zero in $\mathcal{C}_{n}^{\rm Cube}(X)$, being a sum of degenerate $n$-cubes $Q_{n}\rightarrow x_{0}$, so $h^{D}([f]\cdot[g])=h^{D}([f])+h^{D}([g])$. ∎
There is a discrete Hurewicz theorem in dimension $1$. The same strategy in the proof [4, Theorem 4.1] applies in this case.
Theorem 5.4.
For any finite poset $X$, the map
$h^{D}:\pi^{D}_{1}(X,x_{0})\rightarrow{H}^{\rm Cube}_{1}(X)$
is surjective with ${\rm Ker}h^{D}=[\pi^{D}_{1}(X,x_{0}),\pi^{D}_{1}(X,x_{0})]$.
Let us recall the notion of classical Hurewicz map [6]. Thinking of $\pi_{n}(X,x_{0})$ as homotopy classes of maps $f:(L^{n},\partial L^{n})\rightarrow(X,x_{0})$, the Hurewicz map $h:\pi_{n}(X,x_{0})\rightarrow H_{n}^{\rm Sing}(X,x_{0})$ is defined by $h([f])=f_{*}(\beta)$, where $\beta$ is a fixed generator of $H_{n}^{\rm Sing}(L^{n},\partial L^{n})=\mathbb{Z}$ and $f_{*}:H_{n}^{\rm Sing}(L^{n},\partial L^{n})\rightarrow H_{n}^{\rm Sing}(X,x_{0})$ is induced by $f$.
Theorem 5.5.
For any $n\geq 0$, the following diagram commutes.
${\pi_{n}^{D}(X,x_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h^{D}}$$\scriptstyle{\theta}$${H_{n}^{\rm Cube}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi}$${H^{\rm Simpl}_{n}(\mathcal{K}(X))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\eta_{*}}$${\pi_{n}(X,x_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h}$${H^{\rm Sing}_{n}(|\mathcal{K}(X)|).}$
Proof.
Let $f:(\mathbb{Z}^{n},\partial\mathbb{Z}^{n})\rightarrow(X,x_{0})$ be a map of radius $r$. By computing, we get
$\begin{split}\eta_{*}\psi_{*}h^{D}([f])&=\eta_{*}\psi[\sum_{x\in I^{n}_{2r-1}}(T_{r}f)^{x}]\\ &=\eta_{*}\left[\sum_{x\in I^{n}_{2r-1}}\sum_{\tau\in S_{n}}{\rm sign}(\tau)\mathcal{K}\left(\{(T_{r}f)^{x}(p_{\tau}(i))~{}|~{}0\leq i\leq n\}\right)\right]\\ &=\eta_{*}\left[\sum_{x\in I^{n}_{2r-1}}\sum_{\tau\in S_{n}}{\rm sign}(\tau)\mathcal{K}\left(\{T_{r}f(x+p_{\tau}(i))~{}|~{}0\leq i\leq n\}\right)\right]\\ &=\eta_{*}\mathcal{K}(T_{r}f)_{*}\left[\sum_{x\in I^{n}_{2r-1}}\sum_{\tau\in S_{n}}{\rm sign}(\tau)\mathcal{K}\left(\{x+p_{\tau}(i)~{}|~{}0\leq i\leq n\}\right)\right],\end{split}$
where $\left[\sum\limits_{x\in I^{n}_{2r-1}}\sum\limits_{\tau\in S_{n}}{\rm sign}(\tau)\mathcal{K}\left(\{x+p_{\tau}(i)~{}|~{}0\leq i\leq n\}\right)\right]$ is a generator of
$H^{\rm Simpl}_{n}(\mathcal{K}(I^{n}_{2r}),\mathcal{K}(\partial I^{n}_{2r})),$
denoted $\alpha$.
Consider an order-preserving map $P:\mathbb{Z}^{n}\rightarrow\mathbb{Z}^{n}$ given by
$P(x)=x+r.$
Then we have $f(x)=T_{r}f(Px)$ for every $x\in\mathbb{Z}_{\leq r}^{n}$. This says when restricted to $\mathbb{Z}_{\leq r}^{n}$,
$|\mathcal{K}(f)|=|\mathcal{K}(T_{r}f)\circ\mathcal{K}(P)|.$
Denote by $\beta={\lambda_{r}}_{*}|\mathcal{K}(P|\mathbb{Z}_{\leq 2r})|^{-1}_{*}\eta_{*}\alpha$ the fixed generator of $H^{\rm Sing}_{n}(L^{n},\partial L^{n})$. Then
$h\theta([f])$ $=(|\mathcal{K}(f)|\circ\lambda_{r}^{-1})_{*}\eta_{*}(\beta)$
$=|\mathcal{K}(T_{r}f)\circ\mathcal{K}(P|\mathbb{Z}_{\leq 2r})|_{*}\circ{(\lambda_{r}^{-1})}_{*}({(\lambda_{r})}_{*}|\mathcal{K}(P|\mathbb{Z}_{\leq 2r})|^{-1}_{*}\eta_{*}\alpha)$
$=|\mathcal{K}(T_{r}f)|_{*}\eta_{*}\alpha$
$=\eta_{*}\mathcal{K}(T_{r}f)_{*}\alpha.$
Data availability
This study has no associated data.
Declarations
The authors declare that they have no conflict of interest.